What is limiting reagent 3g of h2 reactWith 29g of o2 to produce h2o? calculate Amount of h2o. Produced

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
2H2+O2=2H2O
4g of H2 reacts with 32g of O2 to give 36g of water.
3g of H2 reacts with 24g of O2. But O2 concentration is 29g, therefore it is in excess and H2 is the limiting reagent.
Amount of H2O= 3*36/4=27g.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)